### v4

# First-return statistics in Henyey–Greenstein scattering: Motzkin polynomials and the Cauchy kernel

Abstract
We study first-return statistics for photons undergoing three-dimensional Henyey–Greenstein scattering in a semi-infinite medium. In previous work, we showed that one-dimensional first-passage probabilities expand in Catalan and Motzkin generating functions. Extending to three dimensions requires a Boundary Truncation Factor (BTF) that accounts for the restricted angular phase space imposed by the boundary. Extensive Monte Carlo simulations are used to determine the BTF empirically as a function of scattering order and anisotropy. For moderate anisotropy, the BTF is accurately described by a Cauchy kernel with parameters depending only on the Henyey–Greenstein asymmetry factor. This closed-form expression reproduces Monte Carlo results to 1–2% accuracy over a broad range of scattering orders. At higher anisotropy, systematic deviations from the Cauchy form are observed and can be reduced using a one-parameter generalized kernel. The resulting framework provides a computationally efficient mapping from three-dimensional anisotropic transport to one-dimensional combinatorial first-passage theory.
Keywords: first-passage problem, random walk, Cauchy kernel, Motzkin polynomials, radiative transfer, Henyey–Greenstein scattering
1 Introduction
First-passage problems appear throughout statistical mechanics—in polymer physics, diffusion-limited aggregation, financial mathematics, and queueing theory [1, 2, 3]. A clean example arises in radiative transport: photons entering a scattering medium execute a three-dimensional random walk and may return to the entry boundary. We consider a semi-infinite medium occupying the half-space $z>0$, with the entry boundary at $z=0$; photons enter at normal incidence. The statistics of this first-passage event encode how stochastic motion interacts with geometric constraint.
Throughout, “first-return” means first passage back to $z=0$ from $z>0$ traveling in the z-z direction—the photon must exit, not graze inward.
In our previous work in this journal [6], we established that first-return probabilities in one-dimensional isotropic scattering expand in Catalan numbers, the combinatorial objects counting Dyck paths. The reflectance of a semi-infinite Kubelka–Munk medium [7] admits the generating function representation
$R_{\infty}(S,\chi)=\frac{1}{2}\cdot\frac{S}{S+\chi}\cdot C\left(\frac{S^{2}}{4(S+\chi)^{2}}\right)$ (1)
where $C(x)=\sum_{n=0}^{\infty}C_{n}x^{n}$ is the Catalan generating function [17] and $C_{n}=(2n)!/(n!(n+1)!)$. This is distribution-free: it depends only on the zigzag structure, not step lengths. Forward-peaked scattering introduces “flat” steps—events with no $z$-direction change. The Motzkin extension [14, 15] handles this, giving
$P_{\mathrm{refl}}^{(1D)}(n,r)=M_{n-2}\left(\frac{1-2r}{r}\right)\cdot\left(\frac{r}{2}\right)^{n-1}$ (2)
where $M_{n}(t)$ is the Motzkin polynomial and $r$ the backward-step probability.
The central challenge is extending this framework to three-dimensional anisotropic scattering governed by the Henyey–Greenstein phase function [12]. Direct embedding of Motzkin structure via an effective backscattering coefficient $r_{b}(g)$ fails; as we show below, the mapping requires a Boundary Truncation Factor (BTF) that accounts for geometric constraints at the boundary.
We show that the BTF follows a Cauchy kernel in the scattering order $n$:
$\mathrm{BTF}(n,g)=\frac{A(g)}{1+\left(\frac{n-n_{0}}{m_{x}(g)}\right)^{2}}$ (3)
with parameters expressible in terms of the anisotropy factor $g$ alone: {align} m_x(g) = 4g1-g   $\text$(width)
A(g) = 1 - g(1+g)2   $\text$(amplitude)
n_0 = 2   $\text$(peak location) The simple integer coefficients suggest underlying geometric structure; however, a first-principles derivation of the Cauchy form remains an open problem (Section 7.1).
The theory applies for $g\lesssim 2/3$ and $n\geq 2$; above $g\approx 2/3$, deviations grow gradually but can be corrected by a modified Cauchy kernel with a shape parameter (Section 6).
Practical scope of the validity range. The constraint $g<2/3$ complements rather than competes with biological tissue optics, where $g\approx 0.9$$0.98$ is typical [21]. Table 1 summarizes representative anisotropy factors for various scattering media. Many industrial and environmental applications fall within the validity range of the present theory; for higher-anisotropy materials, the Cauchy kernel form remains valid but parameters benefit from Monte Carlo calibration (Section 4).
Table 1: Representative anisotropy factors $g$ for various scattering media.
Material $g$ Reference
Within validity range ($g<2/3$)
Isotropic scatterers 0 Definition
Paper / print media 0.40.40.60.6 [22]
Above validity range ($g>2/3$)
Human dermis/epidermis 0.70.70.90.9 [20]
Biological tissue (in vivo) 0.90.90.980.98 [21]
Intralipid phantoms $\sim 0.9$ Standard value
1.1 Relation to previous work
Table 2 summarizes the progression from [6].
Table 2: Progression of results from [6] through the present work.
Aspect Zeller & Cordery (2020) Present work
Dimension 1D 3D $\to$ 1D via BTF
Scattering Isotropic + forward bias Henyey–Greenstein
Combinatorics Catalan $\to$ Motzkin Motzkin + BTF correction
BTF Not needed Cauchy kernel (empirical)
Incidence Normal only Normal only
Validity range $g<2/3$ (Cauchy); $g<0.95$ (modified)
Status Derived Empirical; derivation open
From our 2020 paper, we carry forward: (i) the Motzkin polynomial framework (equation (2)); (ii) the distribution-free character of first-passage combinatorics; and (iii) the connection to classical fluctuation theory [5, 4]. What is new here is: (i) identification of the Cauchy kernel form for the BTF through systematic model selection; (ii) determination of the parameters $m_{x}(g)$ and $A(g)$ from Monte Carlo fitting; and (iii) a modified Cauchy kernel extending validity to high anisotropy.
1.2 Comparison with generalized Kubelka–Munk
Our approach differs from previous 3D extensions of Kubelka–Munk. Sandoval and Kim [9, 10] extended KM through double spherical harmonics ($\mathrm{DP}_{1}$), obtaining an $8\times 8$ system for forward and backward power flow. For isotropic scattering in optically thick media, their generalized KM achieves errors below 15% when $z_{0}\geq 10$ (where $z_{0}$ is the optical thickness).
However, $\mathrm{DP}_{1}$ encounters difficulties for anisotropic scattering. Its basis functions contain only first-order azimuthal harmonics and cannot capture forward-peaked phase functions. At $g=0.8$, Sandoval and Kim found errors exceeding 80% in transmitted power; at higher anisotropy, the approximation gives negative intensities.
The BTF framework inverts the dimensional strategy: instead of enriching 1D equations with 3D coupling, we use 1D combinatorics and correct for boundary truncation. Table 3 compares the two approaches.
Table 3: Comparison of dimensional reduction strategies for semi-infinite media.
Property gKM (Sandoval & Kim) BTF (this work)
Geometry Finite slab Semi-infinite
Strategy 1D $\to$ 3D extension 3D $\to$ 1D reduction
Angular basis $\mathrm{DP}_{1}$ (4 func./hemisphere) HG sampling + Motzkin
Isotropic limit <<15% error for $z_{0}\geq 10$ Recovers Catalan structure
Anisotropic range $g\lesssim 0.5$ (qualitative) $g<2/3$ (<<2% deviation)
High-$g$ behavior Unphysical (negative values) Predictable drift
Incidence Oblique via eigenmode Normal (oblique: future work)
Computational form $8\times 8$ PDE system 1D polynomial evaluation
Practitioner guidance. For semi-infinite media with $g\lesssim 2/3$, BTF+Motzkin is more accurate than $\mathrm{DP}_{1}$: 1D polynomial evaluation achieves sub-2% accuracy versus 15%+ for gKM. For finite slabs or strong absorption, use gKM or Monte Carlo. BTF excels in inverse problems requiring thousands of transport evaluations.
1.3 Paper organization
Section 2 reviews the 1D Motzkin framework [6]. Section 3 introduces the Boundary Truncation Factor and its Cauchy kernel form. Section 4 describes the Monte Carlo procedure. Section 5 presents the computational algorithm. Section 6 presents the modified Cauchy kernel for high-anisotropy scattering ($g>2/3$). Section 7 concludes.
2 One-dimensional theory: Catalan and Motzkin structures
We review results from [6]; see [1, 2] for background.
2.1 Kubelka–Munk reflectance and Catalan numbers
The Kubelka–Munk equations [7, 11] describe one-dimensional radiative transport with isotropic scattering:
[…]
where $I$ and $J$ are forward and backward fluxes, $S$ the scattering coefficient, and $\chi$ the absorption coefficient. The reflectance of a semi-infinite slab is
$R_{\infty}(S,\chi)=\frac{S+\chi}{S}-\sqrt{\left(\frac{S+\chi}{S}\right)^{2}-1}$ (5)
The key result of [6] is that this reflectance expands in Catalan numbers:
$R_{\infty}(S,\chi)=\sum_{n_{p}=1}^{\infty}\frac{C_{n_{p}-1}}{2^{2n_{p}-1}}\left(\frac{S}{S+\chi}\right)^{2n_{p}-1}$ (6)
Photon trajectories form zigzag random walks; first passage occurs at the first backward crossing of $z=0$. The probability at the $n_{p}$-th peak is $P(n_{p})=C_{n_{p}-1}/2^{2n_{p}-1}$, connecting to Spitzer’s identity [3] and Andersen’s equivalence principle [5].
2.2 Motzkin extension for forward scattering
Catalan numbers count Dyck paths—walks restricted to up- and down-steps. Forward scattering introduces flat steps (events with no $z$-direction change), requiring the Motzkin extension [14, 15, 16].
Definition (Motzkin polynomial). The Motzkin polynomial of degree $n$ is
$M_{n}(t)=\sum_{k=0}^{\lfloor n/2\rfloor}T(n,k)\cdot t^{n-2k}$ (7)
where $T(n,k)$ are the Motzkin triangle coefficients [17]:
$T(n,k)=\frac{n!}{(n-2k)!\,k!\,(k+1)!}$ (8)
The one-dimensional first-return probability with forward scattering becomes equation (2), encoding three processes: up-steps (positive $z$-projection), down-steps (negative $z$-projection), and flat steps (zero $z$-projection change).
3 The Boundary Truncation Factor
The Motzkin framework applies to one-dimensional scattering. To use it for three-dimensional Henyey–Greenstein transport, we need a mapping—an effective backscattering coefficient $r_{b}(g,n)$ that makes the 1D formula reproduce 3D first-return probabilities. This section introduces the Boundary Truncation Factor (BTF), the correction that makes this mapping work.
3.1 Physical origin
In bulk scattering without boundaries, Pfeiffer and Chapman [13] proved that the Henyey–Greenstein phase function is closed under successive scattering: after $n$ scattering events, the angular distribution of the scattering cosine $\cos\theta=\bm{\Omega}_{0}\cdot\bm{\Omega}_{n}$ is itself a Henyey–Greenstein distribution with parameter $g^{n}$:
$p_{n}(\cos\theta;g)=P_{\mathrm{HG}}(\cos\theta;g^{n})$ (9)
This closure property—not merely the statement $\langle\cos\theta\rangle=g^{n}$—underlies the dimensional reduction from 3D to 1D.
First-passage problems impose geometric constraints that modify this relationship. The requirement that photons return to their entry boundary restricts the accessible angular phase space: trajectories that wander too far forward are less likely to return. This geometric truncation breaks the closure property, effectively reducing the asymmetry parameter beyond the bulk value:
$g_{\mathrm{eff}}^{\mathrm{(constrained)}}=g^{n}\cdot\mathrm{BTF}(n,g)$ (10)
where $\mathrm{BTF}\leq 1$ represents the multiplicative reduction due to boundary constraints. In the bulk limit (no boundary), $\mathrm{BTF}=1$ and we recover equation (9); at a boundary, the incomplete angular integration yields $\mathrm{BTF}<1$.
Formally, the BTF emerges from boundary-constrained angular integration:
[…]
where the constraint ensures the photon crosses below $z=0$ after exactly $n$ scattering events. These nested integrals become analytically intractable beyond $n=3$.
3.2 Empirical discovery
Since direct evaluation of equation (11) is impractical, we determined BTF empirically. Monte Carlo simulations provide exact first-return probabilities for 3D photon transport. The simulations covered $g\in[0.05,0.95]$ (19 values) and $n\in[2,100]$ (99 values) with $10^{8}$ trajectories per $g$ value and 10 independent runs—approximately $10^{10}$ photon histories ($10^{12}$ scattering events, $\sim$500 CPU-hours).
For each $(g,n)$ pair, we extracted the BTF as the value needed to make the 1D Motzkin formula reproduce the Monte Carlo 3D result:
$P_{\mathrm{3D}}^{\mathrm{(MC)}}(n,g)=P_{\mathrm{1D}}^{\mathrm{(Motzkin)}}(n,r_{b}(g,n))$ (12)
where $r_{b}(g,n)$ depends on BTF through the effective anisotropy.
3.3 The Cauchy kernel
Systematic model selection—starting with high-order Padé approximants (rational polynomial fits) and progressively reducing complexity while monitoring cross-validation error—revealed that the optimal functional form is a Cauchy kernel:
$\mathrm{BTF}(n,g)=\frac{A(g)}{1+\left(\frac{n-2}{m_{x}(g)}\right)^{2}}$ (13)
with parameters: {align} A(g) = 1 - g(1+g)2 = (1-g)(2+g)2   $\text$(amplitude)
m_x(g) = 4g1-g   $\text$(width) The peak location $n_{0}=2$ corresponds to the minimum scattering order for first return. This parameterized form reproduces Monte Carlo-derived BTF values with mean deviation <<2% and cross-validated $R^{2}>0.999$ for $g\leq 2/3$.
Statistical evidence for integer coefficients. The 10 independent Monte Carlo runs allow statistical testing of the numerical coefficients. Fitting across all runs yields $n_{0}=2.2\pm 0.3$ and width prefactor $3.9\pm 0.3$; the integers 2 and 4 both lie within one standard deviation of the fitted values. (The exponent 2 in the denominator emerged from the Padé approximant model selection, not from parameter fitting.) This suggests exact integer structure rather than numerical coincidence.
Limiting behavior. For short paths ($n-2\ll m_{x}$), $\mathrm{BTF}\approx A(g)$ with minimal boundary effects. For long paths ($n-2\gg m_{x}$), $\mathrm{BTF}\to 0$ as boundary truncation dominates. At $g=0$ (isotropic), $A=1$ and $m_{x}=0$, so $\mathrm{BTF}=1$ for $n=2$ and $\mathrm{BTF}=0$ for n>2n>2—reflecting that isotropic scattering requires exactly two steps for first return.
Figure 1 shows the BTF directly as a function of scattering order $n$ for several anisotropy values. The Cauchy profile is evident: a peak at $n=2$ with power-law decay at large $n$.
Figure 1: The Boundary Truncation Factor versus scattering order $n$ for $g=0.1,0.3,0.5,0.7,0.9$. Points: extracted from Monte Carlo. Curves: Cauchy kernel (equation 13). The width $m_{x}(g)$ increases with $g$; the amplitude $A(g)$ decreases. At $g=0.9$, deviations from Cauchy become visible.
4 Monte Carlo procedure
This section details the Monte Carlo simulations from which the BTF was extracted.
4.1 Simulation procedure
We follow the standard Monte Carlo procedure for photon transport [18]. First return occurs when $z$ first becomes negative [19]. Photons are initialized at the origin ($z=0^{+}$) with direction cosine $\mu_{0}=1$. At each scattering event, path length is sampled from $p(c)=e^{-c}$ (unit mean free path) and scattering angle from the Henyey–Greenstein phase function via inverse CDF. The scattering order $n$ is recorded at first occurrence of $z<0$. We simulated $10^{8}$ trajectories per $g$ value with 10 independent runs.
The Henyey–Greenstein phase function is
$P_{\mathrm{HG}}(\mu;g)=\frac{1-g^{2}}{2(1+g^{2}-2g\mu)^{3/2}},\quad-1\leq\mu\leq 1$ (14)
where $g=\langle\mu\rangle\in[0,1)$ is the anisotropy parameter.
4.2 Fit quality
The Cauchy kernel captures the Monte Carlo data with mean deviation <<2% for $g<2/3$. Table 4 shows the root-mean-square error (RMSE) and coefficient of determination ($R^{2}$) at representative $g$ values.
Table 4: Fit quality of the Cauchy BTF versus Monte Carlo.
$g$ RMSE $R^{2}$ Max deviation
0.10 $3.2\times 10^{-4}$ 0.9998 1.1%
0.30 $4.1\times 10^{-4}$ 0.9997 1.4%
0.50 $4.8\times 10^{-4}$ 0.9996 1.8%
[…]
0.80 $8.3\times 10^{-4}$ 0.9985 5.2%
0.90 $2.2\times 10^{-3}$ 0.9941 14.8%
Error bounds. For $g<2/3$, root-mean-square deviation is below 2% across all $n$. For $g>2/3$, errors grow approximately as $\epsilon(g)\approx 0.05\cdot(g-2/3)/(1-g)$, reaching $\sim$15% at $g=0.9$. For $g>0.8$, the modified kernel (Section 6) is recommended.
5 Computational algorithm
We present the complete algorithm for mapping 3D Henyey–Greenstein scattering to 1D Motzkin combinatorics. The algorithm has four steps: (1) determine the angular threshold separating forward from backward scattering; (2) compute the BTF-modified exponent; (3) evaluate the effective backscattering probability; (4) apply the Motzkin formula.
Step 1 (Angular threshold): The threshold $\mu_{b}(g)$ is the direction cosine separating “effective backward” from “effective forward” scattering in the 1D projection. Physically, $\mu_{b}$ partitions the angular phase space so that the 1D Motzkin counting matches the 3D return statistics. Solve
[…]
where $p_{r2}(g)=\int_{-1}^{0}\frac{\mu^{\prime}-1}{\mu^{\prime}}P_{\mathrm{HG}}(\mu^{\prime},g)\,\mathop{}\!\mathrm{d}\mu^{\prime}$ is the exact two-step return probability, $F(\mu;g)$ is the HG cumulative distribution function, and $g^{2}$ appears because after two scattering events the angular distribution has parameter $g^{2}$ (equation 9).
Step 2 (Effective exponent): In bulk scattering, the effective anisotropy after $n$ events is $g^{n}$ (standard convolution). Near a boundary, transverse filtering reduces this—captured by the BTF. We model the boundary-constrained effective exponent as
$\gamma(g,n)=2+\mathrm{BTF}(n,g)(n-2)$ (16)
This interpolates between $\gamma=2$ at $n=2$ (where BTF =A(g)=A(g) is maximal and the constraint is weakest) and $\gamma\to 2$ as $n\to\infty$ (where BTF $\to 0$ and boundary truncation dominates). The baseline $\gamma=2$ reflects the minimum scattering order for first return. This ansatz is empirically motivated: it produces the correct limits and matches Monte Carlo across the parameter space.
Step 3 (Effective backscattering): Substitute into the HG CDF:
$r_{b}(g,n)=F(-\mu_{b}(g);g^{\gamma(g,n)})$ (17)
Step 4 (First-return probability): The 3D first-return probability follows from equation (2) with the effective backscattering:
$P_{\mathrm{3D}}^{(\mathrm{refl})}(g,n)=M_{n-2}\left(\frac{1-2r_{b}}{r_{b}}\right)\cdot\left(\frac{r_{b}}{2}\right)^{n-1}$ (18)
This maps 3D Henyey–Greenstein scattering to 1D Motzkin combinatorics.
Isotropic limit check. At $g=0$: $m_{x}=0$, $A=1$, so BTF =1=1 for $n=2$ and BTF =0=0 for $n>2$. The effective exponent becomes $\gamma=2$ for all $n$. The threshold $\mu_{b}(0)=0$ (hemisphere boundary), giving $r_{b}=1/2$. Substituting into equation (18) with $r_{b}=1/2$ recovers the Catalan first-return probabilities of [6], confirming internal consistency.
6 High-anisotropy extension: Modified Cauchy kernel
The Cauchy BTF derived in Section 3.3 works to 2% for $g<2/3$. Biological tissue has $g\approx 0.9$$0.98$ [20, 21], outside this range. Here we develop a one-parameter extension that reduces errors by roughly half at high $g$.
6.1 Systematic deviations at high $g$
Data from $10^{12}$ scattering events show systematic deviations from Cauchy at high $g$. Figure 2 plots $(P_{\mathrm{MC}}-P_{\mathrm{Cauchy}})/P_{\mathrm{Cauchy}}$ against $(n-2)/m_{x}$: for $g>0.85$, Monte Carlo runs high near the peak and low in the tail—lighter tails than Cauchy.
One possible interpretation: strongly forward-peaked scattering may suppress the long, wandering trajectories that populate the Cauchy tail. Photons that scatter many times at high $g$ tend to propagate forward and are less likely to return (ballistic regime); those that do return must do so in relatively few steps. This remains conjecture; a rigorous derivation is lacking.
Figure 2: Deviations from the Cauchy kernel versus scaled path length. At high $g$, Monte Carlo runs high near the peak and low in the tail—the signature of lighter-than-Cauchy tails.
6.2 Modified Cauchy kernel
Setting $\alpha=1$ in the Cauchy form and letting the exponent vary gives the modified BTF:
$\mathrm{BTF}_{\alpha}(n,g)=\frac{A(g)}{\left[1+\left(\frac{n-2}{m_{x}(g)}\right)^{2}\right]^{(1+\alpha(g))/2}}$ (19)
with $A(g)$ and $m_{x}(g)$ unchanged. Setting $\alpha=1$ recovers the standard BTF; $\alpha>1$ gives faster tail decay. This “generalized Cauchy” form appears in robust statistics [23, 24]. Note that $\alpha$ is a shape parameter, not a Lévy stability index.
6.3 Fitting the shape parameter
We fit $\alpha$ at each $g$ by least squares against Monte Carlo. The results follow
[…]
Table 5 compares fitted and calculated values. The parameter stays close to unity: $\alpha\approx 0.98$ at low $g$, rising to $\alpha\approx 1.2$ at $g=0.95$.
Table 5: Fitted $\alpha$ versus formula (20).
$g$ $m_{x}$ $\alpha$ (fitted) $\alpha$ (formula) $(1+\alpha)/2$
0.10 0.4 0.981 0.979 0.99
0.30 1.7 0.990 0.983 0.99
0.50 4.0 0.990 0.989 0.99
[…]
0.80 16 1.018 1.022 1.01
0.90 36 1.090 1.078 1.04
0.95 76 1.196 1.189 1.09
Figure 3 compares Monte Carlo path length distributions with both Cauchy ($\alpha=1$) and $\alpha$-corrected fits. At low $g$, the curves overlap; at high $g$, the $\alpha$-corrected fit (dashed) tracks the data while Cauchy (solid) systematically deviates.
Figure 3: Path length distributions $P(m_{s}|g)$ for six anisotropy values. Gray points: Monte Carlo. Blue solid: Cauchy kernel ($\alpha=1$). Red dashed: Modified Cauchy kernel. At $g\leq 0.7$, the curves are indistinguishable; at $g\geq 0.85$, the modified kernel captures the lighter tails.
The crossover $\alpha=1$ occurs at $g=2/3$, where the Cauchy BTF begins to deviate.
The formula contains the factor $1/(1-g)=\ell^{*}/\ell$, where $\ell$ is the scattering mean free path and $\ell^{*}=\ell/(1-g)$ is the transport mean free path. This ratio appears throughout radiative transfer theory as the natural length scale for direction randomization [8]. Its emergence in the shape parameter suggests a connection to Chandrasekhar’s similarity principle: transport properties depend on $g$ primarily through $\ell^{*}$, not $\ell$ and $g$ separately. Figure 4 plots $\alpha$ against this variable; $\alpha=1$ occurs at $\ell^{*}/\ell=3$.
Figure 4: Shape parameter $\alpha$ versus transport mean free path ratio $\ell^{*}/\ell=1/(1-g)$. The Cauchy case $\alpha=1$ occurs at $\ell^{*}/\ell=3$.
We stress that equation (20) is an empirical fit, not a derivation. It minimizes squared error across all $n$ and captures the trend; the coefficient 0.033 is a fitted constant without theoretical explanation.
6.4 Practical recommendations
Guidelines for implementation:
  • $g<0.85$: use the standard Cauchy BTF. Errors remain under 2%, and the formula is simpler.
  • $0.85\leq g\leq 0.95$: use the modified form with $\alpha$ from equation (20). RMSE improves by $\sim$45%.
  • $g>0.95$: the formula has not been validated beyond $g=0.95$. Monte Carlo calibration is recommended.
For tissue optics ($g\approx 0.9$$0.95$), the modified kernel extends the analytical BTF into the regime of primary interest.
7 Conclusions
The central empirical result is that the Boundary Truncation Factor for 3D Henyey–Greenstein scattering takes a Cauchy kernel form, with parameters $n_{0}=2$, $m_{x}=4g/(1-g)$, and A(g)=1g(1+g)/2A(g)=1-g(1+g)/2. Monte Carlo simulations with $10^{12}$ scattering events confirm this to 1–2% accuracy for $g<2/3$. The parsimony is remarkable: just five parameters capture the entire parameter space (n=2n=2100100, g=0g=00.950.95).
The BTF corrects for transverse diffusion, which reduces return probability at large $n$. The Cauchy form emerges robustly from Monte Carlo. Why this form, and not another, remains open.
For high anisotropy ($g>2/3$), a modified Cauchy kernel with shape parameter $\alpha(g)$ extends the theory to $g\approx 0.95$, covering biological tissue. The high-$g$ extension is phenomenological but constrained by Monte Carlo data—not a free fit.
In practice, the framework replaces 3D Monte Carlo with 1D polynomial evaluation, useful for inverse problems requiring many transport calculations [20, 22].
7.1 Open problems
Several points lack theoretical foundation:
(1) Why Cauchy? The Cauchy kernel emerges empirically but is not derived from first principles. The geometric mechanism—transverse filtering—is clear, but why it produces precisely the Cauchy form (rather than Gaussian, exponential, or other) is unexplained.
(2) Amplitude formula. The expression A(g)=1g(1+g)/2A(g)=1-g(1+g)/2 comes from fitting. A derivation from hemisphere geometry and flux balance would be more satisfying.
(3) Width formula. The factor 4 in $m_{x}=4g/(1-g)$ may decompose as $2\times 2$: one factor from $\langle c^{2}\rangle/\langle c\rangle^{2}=2$ (exponential path lengths), one from first-passage geometry. The geometric factor awaits rigorous analysis.
(4) Shape parameter formula. The empirical fit $\alpha(g)=1+0.033(g-2/3)/(1-g)$ works but is unexplained. The coefficient 0.033 is a fitted constant without evident geometric meaning, unlike the integer coefficients in the Cauchy kernel.
The integer coefficients (2, 4) are not rounded approximations: fitted values across 10 independent Monte Carlo runs are consistent with exactly 2 and 4 within statistical uncertainty. This, combined with the Cauchy kernel form, suggests a clean derivation may exist. We leave this as a challenge for future work.
(5) Oblique incidence. The present work addresses normal incidence only. Extending the Motzkin framework to oblique incidence—where transverse activity $n_{xy}$ couples to the return statistics—requires higher-order Motzkin polynomials, making parameter calculation considerably more challenging. This is an important next step requiring both theoretical development and Monte Carlo validation.
Acknowledgements
CZ thanks Dr. Florence Zeller for discussions on Chebyshev polynomials and Motzkin structures, Professor Arnold Kim for encouragement and guidance, and Professor Michel Talagrand for perspective on the difficulty of proving the Cauchy kernel rigorously.
Appendix A Notation summary
Table 6 summarizes the notation.
Table 6: Principal notation.
Symbol Definition Notes
Scattering parameters
$g$ $\langle\cos\theta\rangle$ Anisotropy factor, $g\in[0,1)$
$\mu$ $\cos\theta$ Direction cosine
$P_{\mathrm{HG}}(\mu;g)$ equation (14) Henyey–Greenstein phase function
$S$ scattering coefficient Units: inverse length
$\chi$ absorption coefficient Units: inverse length
Random walk quantities
$n$ scattering order Number of scattering events
$m_{s}$ path length $\equiv n$; standard MC terminology, used in figures
$n_{p}$ peak count Direction reversals
$n_{f}$ forward-flat count Forward-preserving flat steps
$c$ step length Distance between successive scattering events
Step probabilities
$r_{b}$ backscattering probability 3D $\to$ 1D mapped parameter
$r_{f}$ forward-flat probability Forward-preserving step
$r$ backward-step probability 1D model parameter
BTF parameters
BTF Boundary Truncation Factor Equation (3)
$A(g)$ amplitude 1g(1+g)/21-g(1+g)/2
$m_{x}(g)$ width parameter 4g/(1g)4g/(1-g)
$n_{0}$ peak location =2=2 (minimum scattering order for return)
$\alpha(g)$ shape parameter 1+0.033(g2/3)/(1g)1+0.033(g-2/3)/(1-g); equation (20)
Algorithm quantities
$\mu_{b}(g)$ angular threshold Separates effective forward/backward; equation (15)
$\gamma(g,n)$ effective exponent $2+\mathrm{BTF}(n,g)(n-2)$; equation (16)
$F(\mu;g)$ HG cumulative distribution $\int_{-1}^{\mu}P_{\mathrm{HG}}(\mu^{\prime};g)\,\mathop{}\!\mathrm{d}\mu^{\prime}$
$p_{r2}(g)$ two-step return probability Exact result for $n=2$
Combinatorial objects
$C_{n}$ Catalan number $(2n)!/(n!(n+1)!)$
$M_{n}(t)$ Motzkin polynomial Equation (7)
$T(n,k)$ Motzkin triangle coefficient n!/((n2k)!k!(k+1)!)n!/((n-2k)!\,k!\,(k+1)!)
AI Disclosure
An AI-assisted language model (Claude Opus 4.5, Anthropic, 2025) was used during manuscript preparation to assist in checking and verifying portions of the authors’ own mathematical derivations and to improve clarity of presentation. The AI tool did not generate original results, proofs, or derivations. All mathematics, analyses, and conclusions were independently derived, reviewed, and validated by the authors, who take full responsibility for the content of the manuscript.